# Matryoshka: Non-Exclusive Memory Tiering via Transactional Page Migration

### v1

@@ Abstract
We demonstrate that page migration-based exclusive memory tiering suffers significant performance degradation when fast memory is under pressure. In this paper, we propose non-exclusive memory tiering, a page management strategy that retains a copy of pages recently promoted from slow memory to fast memory to mitigate memory thrashing. To enable non-exclusive memory tiering, we develop Matryoshka, a new page management mechanism for Linux that features transactional page migration and page shadowing. Matryoshka helps remove page migration off the critical path of program execution and makes migration completely asynchronous. Evaluations with carefully crafted micro-benchmarks and real-world applications show that Matryoshka is able to achieve up to 6x performance improvement over the state-of-the-art transparent page placement (TPP) approach in Linux when under memory pressure. We also compare Matryoshka with a recently proposed access sampling-based page migration approach and demonstrate Matryoshka’s strengths and potential weaknesses in various scenarios. Through the evaluation of Matryoshka, we discover a serious issue facing all tested approaches, unfortunately including Matryoshka, and call for further research on tiered memory-aware memory allocation.
@@ 1 Introduction
As new memory devices, such as high bandwidth memory (HBM) [4, 27], DRAM, persistent memory [7, 36], Compute Express Link (CXL)-based memory [1, 41, 34], and storage-class memory [52, 50] continue to emerge, future computer systems are anticipated to feature multiple tiers of memory with distinct characteristics, such as speed, size, power and cost. Tiered memory management aims to leverage the strength of each memory tier to optimize the overall data access latency and bandwidth. Central to tiered memory management is page management within operating systems (OS), including page allocation, placement, and migration. Efficient page management in the OS is crucial for optimizing memory utilization and performance while maintaining transparency for user applications.
@@
Traditionally, the memory hierarchy consists of storage media with at least one order of magnitude difference in performance. For example, in the two-level memory hierarchy assumed by commercial operating systems for decades, DRAM and disks differ in latency, bandwidth, and capacity by 2-3 orders of magnitude. Therefore, the sole goal of page management is to keep hot pages in, and maximize the hit rate of the “performance” tier (DRAM) and migrate (evict) cold pages to the “capacity” tier (disk) when needed. As new memory devices emerge, the performance gap in the memory hierarchy narrows. Evaluations on Intel’s Optane persistent memory  [53] and CXL memory [47] reveal that these new memory technologies are able to achieve comparable performance to DRAM in both latency and bandwidth, within a range of 2-3x. As a result, the assumption on the performance gap, which has guided the design of OS page management for decades, may not hold. It is no longer beneficial to promote a hot page to the performance tier if the migration cost is too high.
@@
Furthermore, unlike disks which must be accessed through the file system as a block device, new memory devices are byte-addressable and can be directly accessed by the processor via ordinary load and store instructions. Therefore, for a warm page on the capacity tier, accessing the page directly and avoiding migration could be a better choice. Most importantly, while the performance of tiered memory remains hierarchical, the hardware is no longer hierarchical. Both the Optane persistent memory and CXL memory appear to the processor as CPUless memory nodes, enabling the OS to manage them as ordinary DRAM nodes.
@@
These unique challenges facing emerging tiered memory systems have inspired research on improving page management in the OS. Much focus has been on expediting page migrations between memory tiers. Nimble [51] improves page migration by utilizing transparent huge pages (THP), multi-threaded migration of a page, and concurrent migration of multiple pages. Transparent page placement (TPP) [41] extends the existing NUMA balancing scheme in Linux to support asynchronous page demotion and synchronous page promotion between fast and slow memory. Memtis [34] and TMTS [22] use hardware performance counters to mitigate the overhead of page access tracking and use background threads to periodically and asynchronously promote pages.
@@
Second, there is a lack of an efficient page migration mechanism to support tiered memory management. As future memory tiers are expected to be addressable by the CPU, page migrations are similar to serving minor page faults and involve three steps: 1) ummap a page from the page table; 2) copy page content to a different tier; 3) remap the page on the page table, pointing to the new memory address. Regardless whether page migration is done synchronously upon accessing a hot page in the slower capacity tier or asynchronously in the background, the 3-step migration process is expensive. During the migration, an unmapped page cannot be accessed by user programs. If page migration is done frequently, e.g., due to memory thrashing, user-perceived bandwidth, including accesses to the migrating pages, is significantly lower (up to 95% lower) than the peak memory bandwidth [51].
@@
We have developed Matryoshka, a new page management mechanism for tired memory that integrates non-exclusive memory tiering and transactional page migration. Matryoshka safeguards page allocation to prevent out-of-memory (OOM) errors due to page shadowing. When the capacity tier is under memory pressure, Matryoshka prioritizes the reclamation of shadow pages before evicting ordinary pages. We have implemented a prototype of Matryoshka in Linux and performed a thorough evaluation. Experimental results show that, compared to two representative page management schemes: TPP and Memtis, Matryoshka achieves up to 6x performance improvement over TPP during memory thrashing and consistently outperforms Memtis by as much as 130% when the working set size fits into fast memory. However, we also discover a potential problem that prevents applications from expanding memory allocations beyond fast memory. This finding motivates more research on the memory allocator in the OS to truly embrace tiered memory.
@@ 2 Motivation and Related Work
We introduce the background of page management in tiered memory systems and use TPP [41], a state-of-the-art page placement system designed for CXL-enabled tiered memory, as a motivating example to highlight the main limitations in current page management approaches.
Figure 1: The comparison of memory bandwidth w/ and w/o TPP upon a CXL-based tiered memory: 16 GB DRAM and 16 GB CXL memory. The workload keeps writing memory by following a Zipfian distribution with two working set sizes (WSS): 10 GB and 24 GB. For the 10GB WSS: 6GB of data on DRAM and 4GB on CXL memory, with 10GB of DRAM pre-allocated to other processes (not involved during testing). For the 24GB WSS: 16GB of data on DRAM and 8GB on CXL memory. TPP in progress shows the performance when TPP starts optimizing memory placement (i.e., during the first round of memory accesses); TPP stable displays the performance once TPP has stabilized (after 4 rounds). High spatial locality: Most frequently accessed data is stored on DRAM. Low spatial locality: Frequently accessed data is evenly distributed between DRAM and CXL memory.
Figure 2: Time breakdown during the initial phase of TPP (i.e., TPP in progress): A significant portion of runtime ($\sim$ 34%) is consumed by TPP’s synchronous page migration.
2.1 The Memory Hierarchy
@@
Caching and tiering are two traditional software approaches to manage a memory, or storage hierarchy, consisting of various types of storage media (e.g., CPU caches, DRAM, and hard disks) differing in performance, capacity, and cost. Without loss of generality, we consider a two-layer memory hierarchy with 1) a performance layer (i.e., fast tier), backed with smaller, faster, but more expensive storage media; and 2) a capacity layer (i.e., slow tier) with larger, slower, and cheaper storage media. For caching, data is stored in the capacity layer, and copies of frequently accessed or “hot” data are strategically replicated to the performance layer. For tiering, frequently accessed data is allocated to the performance layer, while less accessed data is relegated to the capacity layer. At any moment, a unique data copy is maintained exclusively.
Caching and tiering share the same goal – to ensure most accesses occur at the performance layer. But, they differ in how they manage data, in the form of pages within an OS. Caching relies on page replacement algorithms to detect hot (or cold) pages and accordingly populate (or evict from) the performance layer while tiering migrates pages between the performance/capacity layers based on the observed page access pattern (hot or cold). Essentially, caching operates in an inclusive page placement mode and retains pages in its original location, merely storing a copy in the performance layer temporarily for fast access. Conversely, tiering operates in an exclusive mode, actively relocating pages across various memory/storage mediums.
@@
Diverse memory/storage devices, such as high bandwidth memory (HBM) [4], CXL-based memory [1], persistent memory (PM) [7], and fast, byte-addressable NVMe SSDs [28], have emerged recently. While they still make a tradeoff between speed, size, and cost, the distinction between them narrows. For example, Intel Optane DC persistent memory (PM), available in a DIMM package on the memory bus enabling programs to directly access data in non-volatile memory from the CPU using load and store instructions, provides an order of magnitude higher capacity than DRAM (e.g., 8x) and offers performance within an order of magnitude of DRAM, e.g., write latency as low as 80 ns and read latency around 170 ns [53]. More recently, CXL, an open, industry-supported interconnect based on PCI Express (PCIe) interface [1], offers a general interface for disaggregating various types of storage devices (e.g., DRAM, PM, and PCIe devices) to the CPU with memory-like, byte-addressable interface (i.e., via the CXL.mem protocol); it allows cache-line granularity access to the connected devices and maintains data coherency and consistency by the underlying hardware. Real-world CXL-based memory offers comparable memory access latency (<2x) and throughput ($\sim$50%) to local DRAM [47].
CXL memory (or PM), with high throughput and low latency, is often treated as an extension of local DRAM, similar to multi-socket Non-Uniform Memory Access (NUMA) nodes. Hence, a tiering based approach is commonly used to manage such a memory hierarchy. Indeed, state-of-the-art tiered memory systems, such as TPP [41], Memtis [34], Nimble [51], and AutoTiering[29], unanimously adopt tiering to exclusively manage data on different memory tiers, i.e., via dynamically migrating pages between fast/slow tiers to ensure efficient access to faster local DRAM.
However, CXL-based tiered memory has its unique characteristics. CXL memory dramatically increases the capacity of local DRAM, accommodating a broader array of applications with data distributed across both fast and slow memory tiers. As data temperature changes over time [41], more frequent data migrations would occur in CXL-based memory tiers. It is especially true when the fast-tier memory is under pressure and incapable of holding all hot data. A high frequency of data migration can cause memory thrashing, hurting overall system performance. As depicted in Figure 1, the involvement of TPP [41] in page migration (i.e., TPP in progress) has a significant, negative impact on the performance of a memory-intensive workload. In comparison, when most of the hot data have been moved to the fast tier (i.e., TPP stable with the 10GB WSS), the workload achieves higher performance than the case without TPP (i.e., No migration). However, if the hot data is larger than the size of the fast tier (the 24GB WSS case), memory thrashing happens, leading to extremely poor performance (i.e., TPP stable with the 24GB WSS). Furthermore, given that CXL-based lower-tier memory can substantially exceed the size of local DRAM [47], the advantage of storing data exclusively for the sake of efficient memory usage becomes less compelling. Instead, more focus should be placed on reducing the overhead when frequent page migration occurs.
Insight 1: In tiered memory systems, adopting an exclusive, NUMA-like approach for data management can be less effective, particularly when the fast-tier memory cannot accommodate all the hot data.
@@ 2.2 Page Management
Page management has been widely studied to optimize data placement among tiered memory/storage. An effective page management scheme must tackle two primary challenges: 1) transparent and accurate identification of the hotness of pages with minimal overhead used for page placement; and 2) efficient migration of pages between fast and slow tiers without adversely impacting application-level performance.
@@
Tracking memory access can be conducted by software (via the kernel) and/or with hardware assistance. Specifically, the kernel can keep track of page accesses via page faults [2, 41, 29], scanning page tables [17, 2, 40, 51, 14], or both. Capturing each memory access for precise tracking can be prohibitively slow. For example, the page-fault-based tracking approach traps each memory access to selected pages (i.e., whose page table entry permissions are set to no access) via hint page fault. Thus, it allows the kernel to obtain the accurate recency and frequency of each page. However, page faults reside on the critical path of memory accesses, incurring high latency. On the other hand, the scanning-based approach periodically scans page table entries (PTE) and relies on the reference bit to determine which page has been accessed between scanning intervals. However, it monitors page accesses at a coarser granularity, i.e., a longer scanning interval fails to differentiate the frequency and recency of pages, while a short interval incurs high CPU overhead (up to 70% [34]).
To reduce such overhead, Linux adopts a lazy scanning mechanism, which defers the scanning until memory is tight, e.g., when the fast-tier memory is under pressure, and the swapping mechanism is invoked for page demotion (e.g., via kswapd). In addition, the kernel maintains two separate LRU lists for each memory tier: an active list to store hot pages and an inactive list to store cold pages. Pages can be moved between these two lists during (lazy) scanning. The kernel uses two bits (i.e., PG_referenced and PG_active) to maintain the frequency of each page – a cold page, after two accesses (detected by two scans), will be moved to the active list. Due to lazy scanning, it takes a long time to detect hot pages – i.e., lack of recency. However, timely detecting hot pages for the slow tier is critical. TPP [41] tackles this problem by tracking page accesses of the slow tier using hint page faults. However, we observed that TPP’s synchronous page migration approach unexpectedly amplifies its page-fault overhead, e.g., up to 15 page faults were involved in migrating a single detected hot page from the slow to the fast tiers. This is manifested in Figure 2 – a significant portion of time is consumed by “non-migration page faults” (more details in 3.1).
@@
Accurate and lightweight memory access tracking can be achieved with hardware support, e.g., by adding a PTE count field that records the number of memory accesses (by hardware) for the associated page [42]. However, the proposed approach could be too invasive for mainstream architectures (e.g., x86). In practice, the hardware-assisted sampling approach, via Processor Event-Based Sampling (PEBS) [34, 22], has been recently proposed to obtain page access (virtual address) information from sampled hardware events (e.g., LLC misses or store instructions). However, PEBS-based profiling also requires a careful balance between the frequency of sampling and the accuracy of profiling. We observed that the PEBS-based approach [34], with a sampling rate incurring low CPU overhead, remains coarse-grained and fails to capture many hot pages. Further, the sampling-based approach may not obtain recency of memory access (with a low sampling rate), thus limiting the ability to make timely migration decisions. Last, PEBS is not accessible across CPU vendors [22], making the PEBS-based profiling not generic.
Insight 2: A practical, lightweight memory tracking mechanism that can capture both frequency and recency information of page accesses is highly desired.
Page migration – i.e., the process of promoting hot pages from the slow tier to the fast tier or demoting cold pages from the fast tier to the slow tier – involves a complex procedure: ① The system needs to trap to the kernel (e.g., via page faults); ② The PTEs of selected pages must be locked (to avoid others accessing the pages during migration) and ummapped; ③ A translation lookaside buffer (TLB) shootdown must be issued to each processor (via IPI) that may have cached copies of stale PTEs; ④ The content of the pages are copied between tiers; ⑤ Finally, their corresponding PTEs must be updated and remapped. In addition, page migration can be initiated in a synchronously or asynchronously manner. The triggering mode is also linked with the memory tracking mechanisms.
Typically, scanning-based tracking supports asynchronous, batch-based page promotion/demotion, which redistributes pages among fast/slow tiers based on the periodic updates of data temperature [34]. While batching multiple pages for asynchronous migration could reduce migration overhead, it does not favor temporal locality. To migrate on-demand pages more promptly, page-fault-based, synchronous page migration is often used in both Linux and recent research projects [41, 51]: When a “tagged” page is touched for the first time, page migration is synchronously conducted during handling the hint page fault. Despite being timely, synchronous migration is very costly. First, it resides in the critical path of memory access, slowing down the execution of the process requesting such memory access. Further, due to locking (in ①), others trying to access the same page must enter a busy-wait state until the page becomes available remapped). Last, page migration may fail, e.g., due to insufficient space on the destination memory tier, and the need to synchronously re-try further worsens this situation.
This exactly explains the extremely poor performance caused by TPP, as shown in Figure 1, i.e., the cases in “TPP in progress” and “TPP stable” with the 24 GB WSS. The breakdown of time consumption, in Figure 2, further shows that almost 30% of the execution time was spent in the synchronous page promotion (and related) operations, significantly slowing down the workload.
Insight 3: How can we achieve prompt page migration without delaying/pausing the execution of applications?

### v2

# Nomad: Non-Exclusive Memory Tiering via Transactional Page Migration

@@ Abstract
We demonstrate that page migration-based exclusive memory tiering suffers significant performance degradation when fast memory is under pressure. In this paper, we propose non-exclusive memory tiering, a page management strategy that retains a copy of pages recently promoted from slow memory to fast memory to mitigate memory thrashing. To enable non-exclusive memory tiering, we develop Nomad, a new page management mechanism for Linux that features transactional page migration and page shadowing. Nomad helps remove page migration off the critical path of program execution and makes migration completely asynchronous. Evaluations with carefully crafted micro-benchmarks and real-world applications show that Nomad is able to achieve up to 6x performance improvement over the state-of-the-art transparent page placement (TPP) approach in Linux when under memory pressure. We also compare Nomad with a recently proposed hardware-assisted, access sampling-based page migration approach and demonstrate Nomad’s strengths and potential weaknesses in various scenarios.
@@ 1 Introduction
As new memory devices, such as high bandwidth memory (HBM) [hbm, jun2017hbm], DRAM, persistent memory [optane, p2cache], Compute Express Link (CXL)-based memory [cxl, Maruf_23asplos_tpp, memtis], and storage-class memory [optane_study, xiang2022characterizing] continue to emerge, future computer systems are anticipated to feature multiple tiers of memory with distinct characteristics, such as speed, size, power, and cost. Tiered memory management aims to leverage the strength of each memory tier to optimize the overall data access latency and bandwidth. Central to tiered memory management is page management within operating systems (OS), including page allocation, placement, and migration. Efficient page management in the OS is crucial for optimizing memory utilization and performance while maintaining transparency for user applications.
@@
Traditionally, the memory hierarchy consists of storage media with at least one order of magnitude difference in performance. For example, in the two-level memory hierarchy assumed by commercial operating systems for decades, DRAM and disks differ in latency, bandwidth, and capacity by 2-3 orders of magnitude. Therefore, the sole goal of page management is to keep hot pages in, and maximize the hit rate of the “performance” tier (DRAM), and migrate (evict) cold pages to the “capacity” tier (disk) when needed. As new memory devices emerge, the performance gap in the memory hierarchy narrows. Evaluations on Intel’s Optane persistent memory  [pmperf] and CXL memory [Sun2023DemystifyingCM] reveal that these new memory technologies can achieve comparable performance to DRAM in both latency and bandwidth, within a range of 2-3x. As a result, the assumption of the performance gap, which has guided the design of OS page management for decades, may not hold. It is no longer beneficial to promote a hot page to the performance tier if the migration cost is too high.
@@
Furthermore, unlike disks which must be accessed through the file system as a block device, new memory devices are byte-addressable and can be directly accessed by the processor via ordinary load and store instructions. Therefore, for a warm page on the capacity tier, accessing the page directly and avoiding migration to the performance tier could be a better option. Most importantly, while the performance of tiered memory remains hierarchical, the hardware is no longer hierarchical. Both the Optane persistent memory and CXL memory appear to the processor as a CPUless memory node and thus can be used by the OS as ordinary DRAM.
@@
These unique challenges facing emerging tiered memory systems have inspired research on improving page management in the OS. Much focus has been on expediting page migrations between memory tiers. Nimble [Yan_19asplos_nimble] improves page migration by utilizing transparent huge pages (THP), multi-threaded migration of a page, and concurrent migration of multiple pages. Transparent page placement (TPP) [Maruf_23asplos_tpp] extends the existing NUMA balancing scheme in Linux to support asynchronous page demotion and synchronous page promotion between fast and slow memory. Memtis [memtis] and TMTS [Padmapriya_23asplos] use hardware performance counters to mitigate the overhead of page access tracking and use background threads to periodically and asynchronously promote pages.
@@
Second, there is a lack of an efficient page migration mechanism to support tiered memory management. As future memory tiers are expected to be addressable by the CPU, page migrations are similar to serving minor page faults and involve three steps: 1) unmap a page from the page table; 2) copy the page content to a different tier; 3) remap the page on the page table, pointing to the new memory address. Regardless of whether page migration is done synchronously upon accessing a hot page in the slower capacity tier or asynchronously in the background, the 3-step migration process is expensive. During migration, an unmapped page cannot be accessed by user programs. If page migration is done frequently, e.g., due to memory thrashing, user-perceived bandwidth, including accesses to the migrating pages, is significantly lower (up to 95% lower) than the peak memory bandwidth [Yan_19asplos_nimble].
@@
We have developed Nomad, a new page management framework for tired memory that integrates non-exclusive memory tiering and transactional page migration. Nomad safeguards page allocation to prevent out-of-memory (OOM) errors due to page shadowing. When the capacity tier is under memory pressure, Nomad prioritizes the reclamation of shadow pages before evicting ordinary pages. We have implemented a prototype of Nomad in Linux and performed a thorough evaluation on four different platforms, including an FPGA-based CXL prototype, a persistent memory system, and a pre-market, commercial CXL system. Experimental results show that, compared to two representative page management schemes: TPP and Memtis, Nomad achieves up to 6x performance improvement over TPP during memory thrashing and consistently outperforms Memtis by as much as 130% when the working set size fits into fast memory. However, we also discover a potential problem that prevents applications from expanding memory allocations beyond fast memory. This finding motivates more research on the memory allocator in the OS to truly embrace tiered memory.
@@ 2 Motivation and Related Work
We introduce the background of page management in tiered memory systems and use TPP [Maruf_23asplos_tpp], a state-of-the-art page placement system designed for CXL-enabled tiered memory, as a motivating example to highlight the main limitations of current page management approaches.
Figure 1: The comparison of achieved memory bandwidth in a micro-benchmark due to different phases in TPP and a baseline approach that disables page migration. Higher is better performance.
2.1 Memory Tiering
@@
Caching and tiering are two traditional software approaches to manage a memory, or storage hierarchy, consisting of various types of storage media (e.g., CPU caches, DRAM, and hard disks) differing in performance, capacity, and cost. Without loss of generality, we consider a two-level memory hierarchy with 1) a performance tier (i.e., the fast tier), backed with smaller, faster, but more expensive storage media; and 2) a capacity tier (i.e., the slow tier) with larger, slower, and cheaper storage media. For caching, data is stored in the capacity tier, and copies of frequently accessed or “hot” data are strategically replicated to the performance tier. For tiering, new data is first allocated to the performance tier and remains there if it is frequently accessed, while less accessed data may be relegated to the capacity tier when needed. At any moment, data resides exclusively in one of the tiers but not both. Essentially, caching operates in an inclusive page placement mode and retains pages in their original locations, only temporarily storing a copy in the performance tier for fast access. Conversely, tiering operates in an exclusive mode, actively relocating pages across various memory/storage mediums.
@@
Diverse memory/storage devices, such as high bandwidth memory (HBM) [hbm], CXL-based memory [cxl], persistent memory (PM) [optane], and fast, byte-addressable NVMe SSDs [pcieovercxl], have emerged recently. While they still make a tradeoff between speed, size, and cost, the gap between their performance narrows. For example, Intel Optane DC persistent memory (PM), available in a DIMM package on the memory bus enabling programs to directly access data from the CPU using load and store instructions, provides (almost) an order of magnitude higher capacity than DRAM (e.g., 8x) and offers performance within a range of 2-3x of DRAM, e.g., write latency as low as 80 ns and read latency around 170 ns [pmperf]. More recently, compute express link (CXL), an open-standard interconnect technology based on PCI Express (PCIe) [cxl], provides a memory-like, byte-addressable interface (i.e., via the CXL.mem protocol) for connecting diverse memory devices (e.g., DRAM, PM, GPUs, and smartNICs). Real-world CXL memory offers comparable memory access latency (<2x) and throughput ($\sim$50%) to ordinary DRAM [Sun2023DemystifyingCM].
From the perspective of OS memory management, CXL memory or PM appears to be a remote, CPUless memory node, similar to a multi-socket non-uniform memory access (NUMA) node. State-of-the-art tiered memory systems, such as TPP [Maruf_23asplos_tpp], Memtis [memtis], Nimble [Yan_19asplos_nimble], and AutoTiering[Jonghyeon_21fast], all adopt tiering to exclusively manage data on different memory tiers. Unlike the traditional two-level memory hierarchy involving DRAM and disks, in which DRAM acts as a cache for the much larger storage tier, current CXL memory tiering treats CXL memory as an extension of local DRAM. While exclusive memory tiering avoids data redundancy, it necessitates data movement between memory tiers to optimize the performance of data access, i.e., promoting hot data to the fast tier and demoting cold data to the slow tier. Given that all memory tiers are byte-addressable by the CPU and the performance gap between tiers narrows, it remains to be seen whether exclusive tiering is the optimal strategy considering the cost of data movement.
We evaluate the performance of transparent page placement (TPP) [Maruf_23asplos_tpp], a state-of-the-art and the default tiered memory management in Linux. Figure 1 shows the bandwidth of a micro-benchmark that accesses a configurable working set size (WSS) following a Zipfian distribution in a CXL-based tiered memory system. More details of the benchmark and the hardware configurations can be found in Section 4. We compare the performance of TPP while it actively migrates pages between tiers for promotion and demotion (denoted as TPP in progress) and when it has finished page relocation (TPP stable) with that of a baseline that disables page migration (no migration). The baseline does not optimize page placement and directly accesses hot pages from the slow tier. The tiered memory testbed is configured with 16GB fast memory (local DRAM) and 16GB slow memory (remote CXL memory). We vary the WSS to fit in (e.g., 10GB) and exceed (e.g., 24GB) fast memory capacity. Note that the latter requires continuous page migrations between tiers since hot data spills into slow memory. Additionally, we explore two initial data placement strategies in the benchmark. First, the benchmark pre-allocates 10GB of data in fast memory to emulate the existing memory usage from other applications. Frequency-opt is an allocation strategy that places pages according to the descending order of their access frequencies (hotness). Thus, the hottest pages are initially placed in fast memory until the WSS spills into slow memory. In contrast, Random employs a random allocation policy and may place cold pages initially in fast memory.
We have important observations from results in Figure 1. First, page migration in TPP incurs significant degradation in application performance. When WSS fits in fast memory, TPP stable, which has successfully migrated all hot pages to fast memory, achieves more than an order of magnitude higher bandwidth than TPP in progress. Most importantly, no migration is consistently and substantially better than TPP in progress, suggesting that the overhead of page migration outweighs its benefit until the migration is completed. Second, TPP never reaches a stable state and enters memory thrashing when WSS is larger than the capacity of fast memory. Third, page migration is crucial to achieving optimal performance if it is possible to move all hot data to fast memory and the initial placement is sub-optimal, as evidenced by the wide gap between TPP stable and no migration in the 10GB WSS and random placement test.
@@ 2.2 Page Management
In this section, we delve into the design of page management in Linux and analyze its overhead during page migration. We focus our discussions on 1) how to effectively track memory accesses and identify hot pages, and 2) the mechanism to migrate a page between memory tiers.
@@
Tracking memory access can be conducted by software (via the kernel) and/or with hardware assistance. Specifically, the kernel can keep track of page accesses via page faults [autoNUMA, Maruf_23asplos_tpp, Jonghyeon_21fast], scanning page tables [Bergman_22ismm, autoNUMA, Adnan_22hpca, Yan_19asplos_nimble, thermostat], or both. Capturing each memory access for precise tracking can be expensive. Page fault-based tracking traps memory accesses to selected pages (i.e., whose page table entry permissions are set to no access) via hint (minor) page faults. Thus, it allows the kernel to accurately measure the recency and frequency of these pages. However, invoking a page fault on every memory access incurs high overhead on the critical path of program execution. On the other hand, page table (PT) scanning periodically checks the access bit in all page table entries (PTE) to determine recently accessed pages since the last scanning. Compared to page fault-based tracking, which tracks every access on selected pages, PT scanning has to make a tradeoff between scanning overhead and tracking accuracy by choosing an appropriate scanning interval [memtis].
Linux adopts a lazy PT scanning mechanism to track hot pages, which lays the foundation for its tiered memory management. Linux maintains two LRU lists for a memory node: an active list to store hot pages and an inactive list for cold pages. By default, all new pages go to the inactive list and will be promoted to the active list according to two flags, PG_referenced and PG_active, in the per-page struct page. PG_reference is set when the access bit in the corresponding PTE is set upon a PTE check and PG_active is set after PG_reference is set for two consecutive times. A page is promoted to the active list when its PG_active flag is set. For file-backed pages, their accesses are handled by the OS through the file system interface, e.g., read() and write(). Therefore, their two flags are updated each time they are accessed. For anonymous pages, e.g., application memory allocated through malloc, since page accesses are directly handled by the MMU hardware and bypass the OS kernel, the updates to their reference flags and LRU list management are only performed during memory reclamation. Under memory pressure, the swapping daemon kswapd scans the inactive list and the corresponding PTEs to update inactive pages’ flags, and reclaims/swaps out those with PG_reference unset. Additionally, kswapd promotes hot pages (i.e., those with PG_active set) to the active list. This lazy scanning mechanism delays access tracking until it is necessary to reduce the tracking overhead, but undermines tracking accuracy.
TPP [Maruf_23asplos_tpp] leverages Linux’s PT scanning to track hot pages and employs page fault-based tracking to decide whether to promote pages from slow memory. Specifically, TPP sets all pages residing in slow memory (e.g., CXL memory) as inaccessible, and any user access to these pages will trigger a minor page fault, during which TPP decides whether to promote the faulting page. If the faulting page is on the active list, it is migrated (promoted) to the fast tier. Page demotion occurs when fast memory is under pressure and kswapd migrates pages from the inactive list to slow memory.
Figure 2: Time breakdown in the execution of TPP in progress: Synchronous page migration and page fault handling account for a significant portion of the runtime.
@@
Accurate and lightweight memory access tracking can be achieved with hardware support, e.g., by adding a PTE count field in hardware that records the number of memory accesses [7056027]. However, hardware-based tracking can increase the complexity and require extensive hardware changes in mainstream architectures (e.g., x86). In practice, the hardware-assisted sampling, such as via Processor Event-Based Sampling (PEBS) [memtis, Padmapriya_23asplos] on Intel platforms, has been employed to record page access (virtual address) information from sampled hardware events (e.g., LLC misses or store instructions). However, PEBS-based profiling also requires a careful balance between the frequency of sampling and the accuracy of profiling. We observed that the PEBS-based approach [memtis], with a sampling rate optimized for minimizing overhead, remains coarse-grained and fails to capture many hot pages. Further, the sampling-based approach may not accurately measure access recency, thus limiting its ability to make timely migration decisions.
Page migration between memory tiers involves a complex procedure: ① The system must trap to the kernel (e.g., via page faults) to handle migration; ② The PTE of a migrating page must be locked to prevent others from accessing the page during migration and be ummapped from the page table; ③ A translation lookaside buffer (TLB) shootdown must be issued to each processor (via inter-processor interrupts (IPIs)) that may have cached copies of the stale PTE; ④ The content of the page is copied between tiers; ⑤ Finally, the PTE must be remapped to point to the new location. Page migration can be done synchronously or asynchronously. Synchronous migration, e.g., page promotion in TPP, is on-demand triggered by user access to a page and on the critical path of program execution. During migration, the user program is blocked until migration is completed. Asynchronous migration, e.g., page demotion in TPP, is handled by a kernel thread (i.e., kswapd), oftentimes off programs’ critical path, when certain criteria are met. Synchronous migration is costly not only because pages are inaccessible during migration but also may involve a large number of page faults.
Figure 2 shows the run time breakdown of the aforementioned benchmark while TPP is actively relocating pages between the two memory tiers. Since page promotion is synchronous, page fault handling and page content copying (i.e., promotion) are executed on the same CPU as the application thread. Page demotion is done through kswapd and uses a different core. As shown in Figure 2, synchronous promotion together with page fault handling incurs significant overhead on the application core. In contrast, the demotion core remains largely idle and does not present a bottleneck. As will be discussed in Section 3.1, userspace run time can also be prolonged due to repeated minor page faults (as many as 15) to successfully promote one page. This overhead analysis explains the poor performance of TPP observed in Figure 1.